OpenStudy (anonymous):

simplify the following: 1. 3x2 + 4x + 5x2 - 5y 2. (4x - 2y) - (2x + 5y)

OpenStudy (anonymous):

1. 3x2 + 4x + 5x2 - 5y =6+4x+10-5y =4x-5y+16 2. (4x - 2y) - (2x + 5y) =4x-2y-2x-5y =2x-7y